# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v2

@@ 2. $(2)$
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. Then given $f^{(s,0)}(\mathbf{x})$ and $f^{(s,1)}(\mathbf{x})$ ($s=0,1$), we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
Proof:
To recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$ for all $2\leq i\leq b$, it is sufficient to obtain the values of $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$. By Equation 7, the value of $f_{j}(\mathbf{x})$ is known to us for all $j\notin\{j_{1}-1,j_{1},j_{2}-1,j_{2}\}$.
When $j_{1}=1$, we have $p_{j_{1}-1}=0$ and hence $A(\mathbf{x})_{p_{j_{1}-1}}$ is not defined. So we only need to find $A(\mathbf{x})_{p_{j_{2}-1}}$. Let $s\in\{0,1\}$ be such that $j_{2}\equiv s\pmod{2}$. If $s=1$, then $f_{j-1}(\mathbf{x})$ is known for all $j\in K_{s}\setminus\{j_{2}\}$. Let $\delta=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$. Then it follows from the first line of Equation 8 that
$f_{j_{2}-1}(\mathbf{x})\equiv\delta\pmod{2N_{1}}.$
Since $0\leq f_{j_{2}-1}(\mathbf{x})<N_{1}$ and $0\leq\delta<2N_{1}$, we conclude that $f_{j_{2}-1}(\mathbf{x})=\delta$.
If $s=0$, then $j_{1}+1\equiv s\pmod{2}$. Since $j_{2}-j_{1}\geq 2$, we have $j_{1}+1\neq j_{2}$. Then $f_{j-1}(\mathbf{x})$ is known for all $j\in K_{s}\setminus\{j_{1}+1,j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1}+1,j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 8 that
$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$ (9)
$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (10)
Since $0\leq f_{j_{1}}(\mathbf{x}),f_{j_{2}-1}(\mathbf{x})<N_{1}$ and $0\leq\delta^{(0)}<2N_{1}$, we conclude from Equation 9 that $f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})=\delta^{(0)}$. Combining this with Equation 10, we get
$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (11)
$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Since $0\leq j_{1}<j_{2}-1<r\left(A(\mathbf{x})_{1}\right)\leq n/b\leq Q_{1}$, we have $j_{1}\not\equiv j_{2}-1\pmod{Q_{1}}$, which implies that system (11) can be uniquely solved and the unique solution is $\left(f_{j_{1}}(\mathbf{x})\pmod{Q_{1}},f_{j_{2}-1}(\mathbf{x})\pmod{Q_{1}}\right)$. Since $0\leq f_{j_{1}}(\mathbf{x}),f_{j_{2}-1}(\mathbf{x})<N_{1}\leq Q_{1}$, we have $f_{j_{1}}(\mathbf{x})=f_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $f_{j_{2}-1}(\mathbf{x})=f_{j_{2}-1}(\mathbf{x})\pmod{Q_{1}}$.
Now assume $j_{1}>1$. First, suppose $j_{1}\equiv j_{2}\pmod{2}$. Let $\delta^{(0)}=\left(f^{(s,0)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s,1)}(\mathbf{x})-\sum_{j\in K_{s}\setminus\{j_{1},j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$, where $s\in\{0,1\}$ be such that $j_{1}\equiv j_{2}\equiv s\pmod{2}$. Then it follows from Equation 8 that
$f_{j_{1}-1}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
$(j_{1}-1)f_{j_{1}-1}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Similar to previous case, we can uniquely solve $f_{j_{1}-1}(\mathbf{x})$ and $f_{j_{2}-1}(\mathbf{x})$ from this equation system.
Now suppose $j_{1}\equiv s_{1}\pmod{2}$ and $j_{2}\equiv s_{2}\pmod{2}$, where $s_{1},s_{2}\in\{0,1\}$ and $s_{1}\neq s_{2}$. We first show how to calculate $f_{j_{2}-1}(\mathbf{x})$. Since $j_{1}\not\equiv s_{2}\pmod{2}$, it holds that $j_{1}+1\equiv s_{2}\pmod{2}$. Note that $j_{1}+1\neq j_{2}$ since $j_{2}-j_{1}\geq 2$ by assumption. It is clear that $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}$. Let $\delta^{(0)}=\left(f^{(s_{2},0)}(\mathbf{x})-\sum_{j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s_{2},1)}(\mathbf{x})-\sum_{j\in K_{s_{2}}\setminus\{j_{1}+1,j_{2}\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 8 that
$f_{j_{1}}(\mathbf{x})+f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
$j_{1}f_{j_{1}}(\mathbf{x})+(j_{2}-1)f_{j_{2}-1}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Similarly, we can obtain $f_{j_{2}-1}(\mathbf{x})$ by solving this equation system.
It remains to show how to calculate $f_{j_{1}-1}(\mathbf{x})$. If $j_{2}=r\left(A(\mathbf{x})_{1}\right)$, then $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{1}}\setminus\{j_{1}\}$. Therefore,
$f_{j_{1}-1}(\mathbf{x})=\left(f^{(s_{1},0)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1}\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}.$
If $j_{2}<r\left(A(\mathbf{x})_{1}\right)$, then $j_{2}+1\equiv j_{1}\pmod{2}$. It is clear that $f_{j-1}(\mathbf{x})$ is known to us for all $j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}$. Let $\delta^{(0)}=\left(f^{(s_{1},0)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}}f_{j-1}(\mathbf{x})\right)\pmod{2N_{1}}$ and $\delta^{(1)}=\left(f^{(s_{1},1)}(\mathbf{x})-\sum_{j\in K_{s_{1}}\setminus\{j_{1},j_{2}+1\}}(j-1)f_{j-1}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 8 that
$f_{j_{1}-1}(\mathbf{x})+f_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{2N_{1}},$
$(j_{1}-1)f_{j_{1}-1}(\mathbf{x})+j_{2}f_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
We can obtain $f_{j_{1}-1}(\mathbf{x})$ by solving this equation system. ∎
@@
Let $Q_{2}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
@@
$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{2}}$
@@
Since the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are given, the value of $g_{j}(\mathbf{x})$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{2}}$. Then it follows from Equation 12 that
@@
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{2}}.$ (14)
@@
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{2}},$ (15)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{2}}.$
Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{2}$, system (15) can be uniquely solved and the unique solution is
$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{2}},g_{j_{2}}(\mathbf{x})\pmod{Q_{2}}\right).$
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{2}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{2}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{2}}$. ∎
@@
Suppose $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1},Q_{2}$, $f^{(s,\ell)}(\mathbf{x})$, $g^{(s)}(\mathbf{x})$ (where $s,\ell\in\{0,1\}$), $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b_{s,0}<2N_{1}$ ($s=0,1$), $0\leq b_{s,1}<Q_{1}$ ($s=0,1$), $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{2}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f^{(s,\ell)}(\mathbf{x})=b_{s,\ell}\text{ for }s,\ell\in\{0,1\},\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $7\log n+(15b-5)\log\log n+O_{b}(1)$.
@@
Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. By Lemma IV.2, there are at least $2^{n-1}$ $7$-regular sequences when $n\geq 9$. It then follows from the pigeonhole principle that there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot(2N_{1})^{2}\cdot(Q_{1})^{2}\cdot(2N_{2})\cdot Q_{2}\cdot(N_{3})^{2}},$
@@
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+2\log Q_{1}+\log Q_{2}+2\log(2N_{1})+\log(2N_{2})+2\log N_{3}+1$
$=4\log n+2\log Q_{1}+\log Q_{2}+(15b-5)\log\log n+O_{b}(1)$
@@
The last inequality follows from the fact $n/b\leq Q_{1},Q_{2}<2n/b$, which is guaranteed by Lemma II.6. ∎
@@
Our first goal is to recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$. Note that if $j=1$, then $p_{j-1}=0$. So we only need to focus on $\mathbf{x}_{I_{j}^{L}}$ for $j\geq 2$. For $s=0,1$, let
$K_{s}=\left\{2\leq j\leq r\left(U(\mathbf{x})_{1}\right):j\equiv s\pmod{2}\right\}.$
Recall that $\left|I_{j}^{L}\right|\leq b-1$. So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N_{1}-1]$, where $N_{1}=q^{b-1}$. Let $Q_{1}\geq\max\{n/b,N_{1}\}$ be the smallest prime. For $s=0,1$, Define
$f^{(s,0)}(\mathbf{x})=\sum_{j\in K_{s}}\mathbf{x}_{I_{j}^{L}}\pmod{2N_{1}},$ (18)
$f^{(s,1)}(\mathbf{x})=\sum_{j\in K_{s}}(j-1)\mathbf{x}_{I_{j}^{L}}\pmod{Q_{1}}.$

### v3

@@ 2. $(2)$
Suppose that $A\left(\mathbf{x}^{\prime}\right)_{1}$ is obtained from $A(\mathbf{x})_{1}$ by deleting one symbol in $A(\mathbf{x})_{1,I_{j_{1}}}$ and one symbol in $A(\mathbf{x})_{1,I_{j_{2}}}$, where $1\leq j_{1}<j_{2}\leq r\left(A(\mathbf{x})_{1}\right)$ and $j_{2}-j_{1}\geq 2$. Suppose that $I_{j_{1}}=\left[p_{j_{1}-1}+1,p_{j_{1}}\right]$ and $I_{j_{2}}=\left[p_{j_{2}-1}+1,p_{j_{2}}\right]$ are known. Then given $f(\mathbf{x})$, we can recover $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ from $\mathbf{x}^{\prime}$, for all $2\leq i\leq b$.
@@
Let $Q_{1}\geq\max\{N_{2},n/b\}$ be the smallest prime. Define
@@
$g^{(1)}\left(\mathbf{x}\right)=\sum_{j=1}^{r\left(A(\mathbf{x})_{1}\right)}jg_{j}(\mathbf{x})\pmod{Q_{1}}$
@@
Since the values of $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$ are given, the value of $g_{j}(\mathbf{x})$ is known for all $j\neq j_{1},j_{2}$. Let $\delta^{(0)}=\left(g^{(0)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}g_{j}(\mathbf{x})\right)\pmod{2N_{2}}$ and $\delta^{(1)}=\left(g^{(1)}\left(\mathbf{x}\right)-\sum_{j\neq j_{1},j_{2}}jg_{j}(\mathbf{x})\right)\pmod{Q_{1}}$. Then it follows from Equation 9 that
@@
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$ (11)
@@
$g_{j_{1}}(\mathbf{x})+g_{j_{2}}(\mathbf{x})\equiv\delta^{(0)}\pmod{Q_{1}},$ (12)
$j_{1}g_{j_{1}}(\mathbf{x})+j_{2}g_{j_{2}}(\mathbf{x})\equiv\delta^{(1)}\pmod{Q_{1}}.$
Since $1\leq j_{1}<j_{2}\leq n/b\leq Q_{1}$, system (12) can be uniquely solved and the unique solution is
$\left(g_{j_{1}}(\mathbf{x})\pmod{Q_{1}},g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}\right).$
Since $0\leq g_{j_{1}}(\mathbf{x}),g_{j_{2}}(\mathbf{x})<N_{2}\leq Q_{1}$, we have $g_{j_{1}}(\mathbf{x})=g_{j_{1}}(\mathbf{x})\pmod{Q_{1}}$ and $g_{j_{2}}(\mathbf{x})=g_{j_{2}}(\mathbf{x})\pmod{Q_{1}}$. ∎
@@
Suppose $n>2b$ and $b>1$. Let $N_{1},N_{2},N_{3},Q_{1}$, $f(\mathbf{x})$, $g^{(0)}(\mathbf{x})$, $g^{(1)}(\mathbf{x})$, $h_{e}(\mathbf{x})$ and $h_{o}(\mathbf{x})$ be defined as above. For all $0\leq a<2^{4\log n+10\log\log n+O(1)}$, $0\leq b<N_{1}$, $0\leq c_{0}<2N_{2}$, $0\leq c_{1}<Q_{1}$ and $0\leq d_{1},d_{2}<N_{3}$, define the code $\mathcal{C}_{1}$ as
$\mathcal{C}_{1}\triangleq\left\{\mathbf{x}\in\Sigma_{2}^{n}:\begin{array}[]{c}A(\mathbf{x})_{1}\text{ is }7\text{-regular},\\ \eta(A(\mathbf{x})_{1})=a,\\ f(\mathbf{x})=b,\\ g^{(0)}(\mathbf{x})=c_{0},g^{(1)}(\mathbf{x})=c_{1},\\ h_{e}(\mathbf{x})=d_{1},h_{o}(\mathbf{x})=d_{2}\end{array}\right\}$
Then $\mathcal{C}_{1}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, when $n\geq 9$, there is some $a$, $b_{s,\ell}$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that the redundancy of $\mathcal{C}_{1}$ is at most $6\log n+(15b-5)\log\log n+O_{b}(1)$.
@@
Let $N_{0}=2^{4\log n+10\log\log n+O(1)}$. By Lemma IV.2, there are at least $2^{n-1}$ $7$-regular sequences when $n\geq 9$. It then follows from the pigeonhole principle that there is some $a$, $b$, $c_{0}$, $c_{1}$, $d_{1}$ and $d_{2}$, such that
$\left|\mathcal{C}_{1}\right|\geq\frac{2^{n-1}}{N_{0}\cdot N_{1}\cdot(2N_{2})\cdot Q_{1}\cdot(N_{3})^{2}},$
@@
$\rho\left(\mathcal{C}_{1}\right)$ $\leq\log N_{0}+\log N_{1}+\log(2N_{2})+\log Q_{1}+2\log N_{3}+1$
$=5\log n+\log Q_{1}+(15b-5)\log\log n+O_{b}(1)$
@@
The last inequality follows from the fact $n/b\leq Q_{1}<2n/b$, which is guaranteed by Lemma II.6. ∎
IV-C Another construction when $b=2$
Recall that in Section IV-A, the decoding of $A(\mathbf{x})_{i,\left[p_{j_{1}-1},p_{j_{1}}\right]}$ and $A(\mathbf{x})_{i,\left[p_{j_{2}-1},p_{j_{2}}\right]}$ is conducted in two steps: 1) recovering $A(\mathbf{x})_{i,p_{j_{1}-1}}$ and $A(\mathbf{x})_{i,p_{j_{2}-1}}$; 2) recovering $A(\mathbf{x})_{i,I_{j_{1}}}$ and $A(\mathbf{x})_{i,I_{j_{2}}}$. This is because the interval $\left[p_{j-1},p_{j}\right]$ intersects both $I_{j-1}$ and $I_{j}$. A natural question is whether we can handle deletions in the interval $\left[p_{j-1},p_{j}\right]$ in one step, instead of two steps. Next, we show that this is possible when $b=2$.
Recall that in Section IV-A, once the first row $A(\mathbf{x})_{1}$ is recovered, the two deletions in the second row $A(\mathbf{x})_{2}$ are determined in two non-overlapping intervals $\left[p_{j_{1}-1},p_{j_{1}}\right]$ and $\left[p_{j_{2}-1},p_{j_{2}}\right]$ in $[n/2]$ if we assume $p_{j_{2}-1}>p_{j_{1}}$ (otherwise, the two deletions can be corrected using the method in Section IV-B). In Lemmas 21, 22 and 23 in [42], Lu et al proved the follows result, which is a straightforward corollary of Lemma 5 and Observation 2 in [19].
Lemma IV.8
[42, Lemma 23] Let $m$, $P_{1}$ and $P_{2}$ be positive integers, where $P_{1}+P_{2}<m$. Let $P=\min\{P_{1},P_{2}\}$. There is a function $\mathcal{E}_{2}:\Sigma_{2}^{m}\rightarrow\Sigma_{2}^{2\log m+\log P+O(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{m}$, if $\mathbf{y}$ is obtained from $\mathbf{x}$ by deleting two bits at two known non-overlapping intervals of length at most $P_{1}$ and $P_{2}$ respectively, then given $\mathcal{E}_{2}(\mathbf{x})$ and $\mathbf{y}$, one can uniqely and efficently recover $\mathbf{x}$.
Note that $p_{j_{1}}-p_{j_{1}-1},p_{j_{2}}-p_{j_{2}-1}\leq 7\log(n/2)+1$. So, let $P_{1}=P_{2}=7\log(n/2)+1$ and $m=n/2$. Then we can apply the function $\mathcal{E}_{2}$ in Lemma IV.8 to handle interval $\left[p_{j_{1}-1},p_{j_{1}}\right]$ (and $\left[p_{j_{2}-1},p_{j_{2}}\right]$) in one step, introducing only $2\log n+\log\log n+O(1)$ additional redundant bits. Now replacing the six functions $f^{(s,\ell)}$, $g^{(}0)$ and $g^{(}1)$ in the definition of $\mathcal{C}_{1}$ by only one function $\mathcal{E}_{2}$, we have the following corollary.
Corollary IV.1
Let $b=2$ and $n\geq 9$. There is a binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $6\log n+25\log\log n+O(1)$.
@@
Our first goal is to recover $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$. Recall that $\left|I_{j}^{L}\right|\leq b-1$. So we can view $\mathbf{x}_{I_{j}^{L}}$ as an integer in $[0,N-1]$, where $N=q^{b-1}$. Note that $I_{1}^{L}=\emptyset$. For uniform discussion, we also let $\mathbf{x}_{I_{1}^{L}}=0$ and $\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}=\mathbf{x}_{\left[n-b+2,n\right]}$. Now define $L(\mathbf{x})=\mathbf{x}_{I_{1}^{L}}\cdots\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}}\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}\in[0,N-1]^{r\left(U(\mathbf{x})_{1}\right)+2}$. In other word, sequence $L(\mathbf{x})$ is the concatenation of $\mathbf{x}_{I_{1}^{L}},\ldots,\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)}^{L}},\mathbf{x}_{I_{r\left(U(\mathbf{x})_{1}\right)+1}^{L}}$. By Equation 14, the value of $\mathbf{x}_{I_{j}^{L}}$ is known to us for all $j\notin\{j_{1},j_{1}+1,j_{2},j_{2}+1\}$.
Since $j_{2}\geq j_{1}+2$, it follows that $L(\mathbf{x})$ suffered two bursts (of length $2$) of erasures. To determine the values of $\mathbf{x}_{I_{j_{1}}^{L}}$ and $\mathbf{x}_{I_{j_{2}}^{L}}$, we apply Lemma II.7. Let $\mathcal{E}_{1}(\cdot)$ be the function defined in Lemma II.7. Define
$f(\mathbf{x})=\mathcal{E}_{1}\left(L(\mathbf{x})\right).$
Since $r(U(\mathbf{x})_{1})\leq n/b$, the vector $f(\mathbf{x})$ can be viewed as an integer in $\left[0,N_{1}-1\right]$, where $N_{1}=2^{\log(n/b+2)+10\log N+4}=2^{\log n+O_{q,b}(1)}$. Now the following lemma follows from above analysis and Lemma II.7.